OpenStudy (anonymous):

Find all the values of x satisfying the given conditions. f(x)= 5x^3+x^2-80x-4 and f(x)=12 What is the solution set?

OpenStudy (anonymous):

I did 5x^3+x^2-80x-4=12-12 5x^3+x^2-80x-4=0 (5x^3+x^2)-(80x-4)=0

OpenStudy (anonymous):

if you subtract 12 you should get \(5x^3+x^2-80x-16=0\)

OpenStudy (anonymous):

How is it -16?

OpenStudy (anonymous):

you were right o:

OpenStudy (anonymous):

@genius12

OpenStudy (anonymous):

You know this is a third degree function with a positive leading coefficient. So the graph goes from the third quadrant to the first quadrant. To find the value of x where f(x) = 12, just equate the two sides. 5x^3 + x^2 - 80x - 4 =12 5x^3 + x^2 - 80x - 16 =0 --> Bring 12 over to the other side. x^2(5x + 1) - 16(5x + 1) = 0 --> 'Distributive Property'. Factored 'x^2' from first two terms and - 16 from the other two. (x^2 - 16)(5x + 1) = 0 (x - 4)(x + 4)(5x + 1) = 0 --> x = 4, -4, -1/5 Now our polynomial is completely factored, which means we can find the solutions. Therefore, the solutions are: x = 4, -4, -1/5 @ikoreanx3